The radical centre of the circles xsquare+ysquare=1 and xsquare +ysquare-2x-4=0 and xsquare+ysquare-3y-3=0 is_

Grade:12

1 Answers

Arun
4 years ago
Dear student
 
radical axis :
S1 – S2 = 0
 
-1 + 2x + 4 = 0
 
x = – 3/2
 
now S1 – S3 = 0
 
- 1 + 3y + 3 = 0
 
y = – 2/3
 
hence radical centre : ( – 3/2, – 2/3)